OpenStudy (anonymous):

Find the equation of the parabola with focus (5, 1) and directrix y = -1.

OpenStudy (anonymous):

ok there you are first lets draw a picture so we see what it looks like kinda

OpenStudy (anonymous):

ok .

OpenStudy (anonymous):



OpenStudy (anonymous):

the vertex is half way between the focus and the directrix that means it is at \((5,0)\) also we see it opens up, so the \(x\) term will be squared not the \(y\) term

OpenStudy (anonymous):

general form is \[4p(y-k)=(x-h)^2\] in your case \(h=5,k=0\)

OpenStudy (anonymous):

\(p\) is the distance between the focus and the vertex, which is 1 so it is \[4y=(x-5)^2\] done

OpenStudy (anonymous):

if you have multiple choice problem and don't see that answer, i could be written as \[y=\frac{1}{4}(x-5)^2\] we can check it if you like

OpenStudy (anonymous):

my choices are : (x-5)^2 = 4y (y-5)^2 = 4x (x-5)^2 = -4y (y-5)^2 = -4x

OpenStudy (anonymous):

\[4y=(x-5)^2\] is the same as \[(x-5)^2=4y\]right?

OpenStudy (anonymous):

yes because a+z = z+a i think .

OpenStudy (anonymous):

can you help me with another one ?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

sure post again